# Rethinking Whole-Body CT Image Interpretation: An Abnormality-Centric Approach

### v1

@@ A.8 Ablation Study on Context Range
A.9 Ablation Study on Different Visual Prompts
Recent studies have revealed that VLMs exhibit varying perceptual capabilities for different forms of visual prompts [8]. Therefore, in the visual-prompted report generation task, we investigated the impact of different visual prompt formats to give insight on optimal design choices in application. Specifically, for each abnormality in our dataset, we simulated four distinct visual prompt types: center cropping, ellipse, contour, and bounding box, as input and evaluated OminiAbnorm-CT’s generated reports respectively. We also take the maximum score among them for each abnormality, representing the theoretical best performance achievable by selecting the optimal visual prompt type for each specific abnormality, depending on its shape, location, and etc. As shown in Table 8, cropping the prompted area consistently underperforms across nearly all metrics (24 out of 27 total measurements), likely due to the loss of crucial contextual information outside the prompted region. The other individual visual prompting methods (ellipse, contour, and bounding box) demonstrate comparable performance. Notably, integrating multiple visual prompting techniques yields a substantial improvement across all metrics compared to the best individual prompt for each measure. This indicates significant performance variation when using different visual prompts for the same abnormality, suggesting that selecting appropriate prompt types based on abnormality morphology and location could be an effective strategy.
Table 8: Comparison of different visual prompt formats in visual prompted report generation task, with OminiAbnorm-CT fixed as baseline. Abbrevations: B = BLEU, RTSc = RaTEScore, BTSc = BERTScore, MTR = METEOR, R = Rouge, RG = RadGraph.
Prompt Type B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
@@ A.9 Ablation Study on Different Visual Prompts
A.10 Generation of Text Queries and Visual Prompts
@@
A.11 Detailed Distribution of Abnormalities in OminiAbnorm-CT-14K
Figure 8: Distribution of OminiAbnorm-CT-14K across 82 anatomical structures and 40 major systems or organs. Darker blue indicates higher sample density. Abd. & Pelv. = Abdominal and Pelvic Cavity; Test., Epid., Scrot. = Testes, Epididymis, Scrotum; Morph. = Morphology; Pos. = Positional; Vent. & Subarach. = Ventricles and Subarachnoid.
Table 9: Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
@@ A.11 Detailed Distribution of Abnormalities in OminiAbnorm-CT-14K
Table 10: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Middle Ear Middle ear fluid or hemorrhage 5 0 0
@@
Table 11: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Table 12: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
@@
Myocardium Myocardial hypertrophy 4 1 1
@@
Table 13: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Heart Coronary arteries Coronary artery myocardial bridge 1 0 0
@@
Table 14: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Gallbladder wall Gallbladder wall thickening 53 11 1
@@
Table 15: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
@@
Table 16: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Table 17: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Table 18: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.

### v2

@@ B.2 Ablation Study on Context Range
Table 7: Impact of multi-slice context range on visual prompted report generation, with the base Qwen2.5-VL-7B model as baseline.
B.3 Ablation Study on Different Visual Prompts
Recent work shows VLMs respond exhibit varying perceptual capabilities for different prompt formats [cai2024vip]. We therefore study visual prompt design for visual prompted report generation. For each abnormality, we simulate four prompts, i.e., center crop, ellipse, contour, and box, and evaluate OmniAbnorm-CT separately. We also consider a per-case oracle that takes the best score among the four, representing the theoretical best performance via selecting the optimal prompt for each specific abnormality, depending on its shape, location, and etc. As shown in Table 8, cropping performs worst on most metrics (24/27), likely due to lost context. Ellipse, contour, and box yield similar performance. The oracle notably improves all metrics over any single prompt, revealing large per-case variance. This suggests that adaptively choosing the visual prompt by morphology and location is beneficial, whereas cropping away context is harmful.
Prompt Type BLEU-1 BLEU-2 BLEU-3 RaTEScore BERTScore METEOR ROUGE-1 ROUGE-L RadGraph
@@ B.3 Ablation Study on Different Visual Prompts
Table 8: Comparison of different visual prompts in visual prompted report generation task, with OmniAbnorm-CT fixed as baseline.
Appendix C Details of AbnormRubric
As introduced in Section 6.2, the computation of AbnormRubric is decomposed into three stages that explicitly mirror the way radiologists assess a report: (i) abnormality detection (recall), (ii) hallucination checking (precision), and (iii) description accuracy of radiological attributions. All three stages are implemented using LLM-based judgments given the ground-truth report and the generated report.
Detection of ground-truth abnormalities (Recall). Given the set of abnormalities in the ground-truth report, we iterate over each and query an LLM to determine whether this abnormality is explicitly present in the generated report. Abnormalities that are judged as present are counted as true positives (TP), while the remaining ground-truth abnormalities are counted as false negatives (FN). The detailed prompt is:
You are a medical imaging report analyst. Your task is to evaluate an AI-generated radiology report (or only a portion/fragment of a report) from two aspects: 1. Whether the report is a VALID radiology report (it makes a clear statement about what abnormalities are present OR absent on the current image). 2. Whether the report identifies a specific abnormality. ABNORMALITY TO CHECK: Target Abnormality: {abnormality_category} Relevant Description in Detailed: {groundtruth_description} AI-GENERATED REPORT TO ANALYZE: {generated_report} EVALUATION CRITERIA: 1. Validity Check: The generated report may contain medically irrelevant content, nonsensical phrases or garbled text. Completely ignore such content and focus only on medically relevant content. The generated report may be blank, or entirely composed of non-medical gibberish, irrelevant text, or insufficient to determine presence/absence of abnormalities. Then ignore the following procedure and your evaluation is complete: is_invalid_report=1, detected=0, description=None. 2. Detection: Report as detected=1 if the target abnormality ({abnormality_category}) is reported in the generated report. Consider synonyms and clinically equivalent terms as matches. Regardless of the description accuracy, such as incorrect location, size measurements, morphological characteristics, density/attenuation, and so on. The generated report may indicate that no abnormalities were detected on the image, for example, I don’t see any relevant abnormalities on the image. Then ignore the following procedure and your evaluation is complete: detected=0. 3. Description Extraction: If detected=1, collect ALL relevant description from the report. Preserve original wording without modifications. If multiple descriptions exist, concatenate them. May include adjacent text if the description cannot be grammatically isolated. REQUIRED OUTPUT FORMAT:
  "is_invalid_report": 0, // 1 or 0
  "detected": 1,      // 1 if detected, 0 if not
  "description": "exact text" // string if detected, null if not
STRICT PROHIBITIONS: NEVER add information not present in the generated report. NEVER paraphrase or summarize original text. DO NOT output any additional text except for a valid json.
Identification of hallucinated abnormalities (Precision). To assess hallucinations, we further instruct the LLM to enumerate any abnormalities that are described in the generated report but do not exist in the ground-truth report. Abnormalities that cannot be matched are counted as false positives (FP). The detailed prompt is:
Your task is to evaluate the hallucinations in an AI-generated radiology report (or only a portion/fragment of a report): 1. Suppose some ground truth abnormalities have been successfully detected by the AI-generated radiology report. 2. Identify any ADDITIONAL abnormalities mentioned in the generated report that are NOT covered by the ground truth list. THE AI-GENERATED REPORT TO ANALYZE: generated_report GROUND TRUTH ABNORMALITIES AND MATCHING RESULTS: groundtruth_abnormalities INSTRUCTIONS: 1. Review matched content: The above shows which ground truth abnormalities were detected and their corresponding text in the generated report. 2. Find additional abnormalities: Identify ALL ADDITIONAL abnormal findings mentioned in the generated report that are NOT already matched to ground truth abnormalities. 3. Exclude already matched content: Do NOT include any abnormalities that are already matched to ground truth abnormalities. 4. Ignore normal findings: Only report abnormalities, not normal findings. 5. Output: ONLY valid JSON with no additional text. NOTE: 1. The generated report may contain medically irrelevant content, nonsensical phrases or garbled text. Completely ignore such content and focus only on medically relevant content. 2. If the generated report is entirely composed of non-medical gibberish, irrelevant text, or is blank, then your evaluation is complete: count = 0 and hallucinated_abnormalities = []. REQUIRED OUTPUT FORMAT:
  "hallucinated_abnormalities": [...],  // list of hallucinated abnormalities
  "count": x  // number of hallucinated abnormalities
If no additional abnormalities found:
  "hallucinated_abnormalities": [],
  "count": 0
Accuracy of radiological attributions. For each detected ground-truth abnormality (TP), we assess how accurately its key radiological attributions are described in the generated report. Based on our abnormality taxonomy, radiologists define a set of clinically important attributions for each abnormality type (e.g., location, morphology, density, size). For a given abnormality, we consider only those attributions that are explicitly mentioned in the ground-truth report and ask the LLM to compare them with the corresponding text in the generated report. For each attribution, the LLM returns a binary judgment of whether the generated description is clinically consistent with the ground-truth description. The detailed prompt is:
You are a medical imaging report analyst. For a single abnormality in the image, the AI model has generated a description. Your task is to compare the AI-generated description with the ground-truth description and evaluate the AI’s accuracy across specific attributes that will be provided. EVALUATION INSTRUCTIONS: INPUT DATA: Abnormality: {abnormality_category} Ground-Truth (GT) Description: {gt_description} AI-Generated (Pred) Description: {pred_description} Attributes: {attribution_ls}
1. Detection of Attribute Mention: Evaluate ONLY the pre-defined attributes provided in the list. For each attribute, check if it is mentioned in the GT description. For attributes present in the GT description, check if they are also mentioned in the Pred description. 2. Comparison: For attributes present in BOTH GT and Pred: Extract the exact phrasing from both descriptions. Determine if the descriptions are clinically equivalent (1) or not (0). Consider synonyms and clinically equivalent terms as matches. Provide brief reasoning for your judgment. For attributes ONLY present in GT: Mark as not clinically equivalent (0). 3. Quantitative Data Handling: If GT contains quantitative data (e.g., measurements with units): Extract the numerical value and unit from GT. Check if Pred provides quantitative data for the same attribute. If Pred lacks quantitative data: mark as non-equivalent (0). If Pred provides data: calculate percentage error after unit conversion. Error $\leq 20\%$: equivalent (1); Error $20\%$: non-equivalent (0). In the reasoning field, include the brief calculation process. 4. Special Cases: If Pred provides quantitative data but GT does not: mark as non-equivalent (0). For ambiguous cases where clinical equivalence cannot be determined: mark as uncertain (1-1). 5. Strict Constraints: Base ALL judgments strictly on the provided GT and Pred texts. NO additions, deletions, modifications, or external inferences. Preserve exact wording when extracting descriptions. Clinical relevance overrides literal text matching: Ignore differences in writing style, terminology preferences, and phrasing variations. Consider clinically synonymous terms as equivalent. Focus on clinical semantic meaning rather than exact word matching. Account for standard medical abbreviations and their equivalent expressions. REQUIRED OUTPUT FORMAT (JSON only):
  "attributes": {
    "attribute_name_1": {
      "present_in_gt": 1, // 0 if not present, and the followings can be empty
      "gt_description": "exact text from GT",  // null if not present
      "pred_description": "exact text from Pred",  // null if not present
      "equivalent": 1,  // 0 if not equivalent, -1 if can not be determined
      "reasoning": "brief clinical justification",  // null if not present
      "gt_quantified": 1,  // 0 if not quantified
      "pred_quantified": 1,  // 0 if not quantified
      "error_percentage": 5.2 // null if not both quantified
      or can not be calculated
Appendix D Supplementary Implementation Details
D.1 Supplementary Implementation Details of OmniAbnorm-CT
Training Hyperparameters. We adopt 8 NVIDIA A100-80G GPUs for training, with 1 sample per device and gradient accumulation set to 4. We train OmniAbnorm-CT for 250K iterations in total, and use 5% steps for warming up. We take AdamW [loshchilov2017decoupled] as optimizer with a learning rate of 2e-4. When optimizing the multi-modal language model and the segmentation module jointly, we assign equal weights to the text generation loss and segmentation loss, as defined in Equation 8. We set the rank of LORA to 32.
Training Data. We pad and rescale all images to $\times512$. For training data sampling from four tasks: visual prompted report generation, grounded report generation, text-guided grounded report generation, and general VQA, we control the ratio to be 1:1:1:1. Additionally, in the abnormality grounding task, we maintain a ratio of 6:4 between OmniAbnorm-CT-14K and public lesion segmentation data. For each abnormality description, we use the following prompt to have GPT rewrite three different versions, then randomly select one as the ground truth:
D.2 Supplementary Implementation Details of Baselines
Visual prompted generation. we prompt each abnormality with all types of visual prompts, and take the maximum score. All the methods are prompted with the following template:
Grounded report generation. We integrate LLaVA-Med with 2D segmentation models as baselines. Specifically, we re-implement the MedULS [de2024uls23] with a 2D nnU-Net [nnUNET] and the public lesion segmentation datasets in Table 9, covering all the datasets in the official ULS-23 challenge and 10 additional ones. To our knowledge, this represents the segmentation model with the broadest capability range (in terms of abnormality variety) that can be constructed from currently available public datasets. For BiomedParse [zhao2024foundation] and LiSA [lai2024lisa], we prompt them with ‘Abnormal findings on the CT image’ to derive segmentation results for all the abnormalities on the input CT image. Then, we combine each segmentation model with LLaVA-Med by converting the segmentation results into bounding boxes overlaid on the CT image, and prompt LLaVA-Med to generate the report based on these visual cues:
Text-guided grounded report generation. We use the simulated text queries as prompts for BiomedParse and LiSA to derive segmentation results for the queried abnormalities, which are detailed in Section 4.3. For MedULS, since it doesn’t support text-prompted segmentation, we simply use its unconditioned segmentation results. Similarly, we convert their segmentation results into bounding boxes overlaid on the CT image, and prompt LLaVA-Med to generate the report based on them:
D.3 Runtime and Computation Analysis
Training. We use 4 A100-80G GPUs for training OmniAbnorm-CT. The segmentation module is first pre-trained for 50,000 steps with batch size 16, which takes around 10 hours. Then the Qwen2.5-VL is fine-tuned on the visual-prompted generation task for 20,000 steps, with batch size of 1 and gradient accumulation every 8 steps, which takes around 8 hours. Finally, the VLM and segmentation modules are jointly trained for grounded report generation, with batch size of 1 and gradient accumulation every 8 steps. This final stage takes approximately 48 hours for 100,000 steps.
Inference. All inferences are conducted on 1 A100-80G GPU. We use mixed-precision without quantization. The average inference latency is 0.49 second per sample for visual-prompted generation, 0.98 second per sample for grounded report generation, and 1.01 second per sample for text-guided grounded report generation.
Appendix E Supplementary Details of OmniAbnorm-CT-14K
E.1 Quality Verification
Annotation Quality. We conducted rigorous and comprehensive quality verification on the annotations in OmniAbnorm-CT-14K. For each annotator, we randomly sampled 100 annotated images on the axial plane, 50 on the coronal, and 50 on the sagittal plane. Then a senior radiologist with 12 years of experience assesses the annotation quality on four key metrics: (i) Detection rate measures the percentage of abnormalities successfully identified and annotated by the annotator, without any omission; (ii) Grounding precision evaluates the percentage of grounding annotations that properly encompass the primary regions of the abnormality, while minimizing false positive areas; (iii) Report concordance quantifies the percentage of description annotations that faithfully reflect the linked abnormal findings on the image; (iv) Classification accuracy measures the percentage of abnormalities that are correctly categorized. The verification results demonstrated that: (i) 3 out of 4 annotators achieved perfect scores across all four metrics (100%). (ii) Only one annotator had 10 abnormality annotations where the descriptions were not sufficiently accurate and the category labels were also incorrect, resulting in Report concordance and Classification accuracy of 95%, while maintaining 100% Detection rate and Grounding precision. These results confirm the high quality of annotations in our OmniAbnorm-CT-14K.
Cross-annotator Consistency. We further conducted inter-annotator consistency checks across 4 annotators on 40 random samples. For each sample, we enumerated all annotator pairs and, for each pair, computed the DSC between their segmentation masks and the BLEU-1 score between their abnormality-level report descriptions. Averaging these metrics across all annotator pairs yielded mean pairwise scores of DSC=84.5 for segmentation and BLEU-1=72.2 for finding descriptions, indicating reasonably good agreement among the annotators.
E.2 Mitigation of the Long-Tail Distribution
To mitigate the long-tail distribution in OmniAbnorm-CT-14K, we identified underrepresented organs in our annotated corpus and strategically employed GPT-4o to analyze unlabeled reports. Using the following prompt, we efficiently filtered cases containing abnormal findings related to these underrepresented organs:
E.3 Comparison against Existing Datasets
We compare OmniAbnorm-CT-14K with existing public CT image datasets in Table 9, including those widely used for lesion segmentation, lesion detection, organ segmentation, or report generation. In contrast to the substantial limitations exhibited by these datasets, as detailed the main paper, OmniAbnorm-CT-14Kepresents the first large-scale dataset designed for abnormality grounding and description across multi-plane and whole-body CT images.
Dataset Task Anatomy Plane Report #Category #Image
ULS Bone [de2024uls23] Lesion Seg. Bone Axial
1 151
ULS Pancreas [de2024uls23] Lesion Seg. Pancreas Axial
1 119
MSD Liver [MSD] Lesion Seg. Liver Axial
1 131
MSD Lung [MSD] Lesion Seg. Lung Axial
1 63
MSD Colon [MSD] Lesion Seg. Colon Axial
1 126
MSD Pancreas [MSD] Lesion Seg. Pancreas Axial
1 281
COVID19 [COVID19] Lesion Seg. Lung Axial
1 20
KiTS23 [KiTS23] Lesion Seg. Kidney Axial
2 489
KiPA22 [KiPA22] Lesion Seg. Kidney Axial
2 70
NSCLC [NSCLC] Lesion Seg. Lung Axial
1 85
LIDC IDRI [LIDCIDRI] Lesion Seg. Lung Axial
1 750
LNDb [LNDb] Lesion Seg. Lung Axial
1 236
INSTANCE22 [INSTANCE] Lesion Seg. Brain Axial
1 100
Seg.Rap2023 [SegRap2023] Lesion Seg. Head & Neck Axial
2 120
FUMPE [FUMPE] Lesion Seg. Lung Axial
1 35
RibFrac [ribfracchallenge2025] Lesion Seg. Rib Axial
1 420
Adrenal ACC Ki67 [AdrenalACCKi67] Lesion Seg. Adrenal Gland Axial
1 29
LNQ2023 [dorent2024lnq] Lesion Seg. Lung Axial
1 393
NIH-LN [roth2014new] Lesion Seg. Lung Axial
1 175
CCC-18 [CCC18] Lesion Seg. Chest & Abdomen Axial
404
DeepLesion [yan2018deep] Lesion Det. Whole Body Axial
33K
TotalSegmentator [Totalsegmentator] Organ Seg. Whole Body Axial
1.2K
AbdomenAtlas [abdomenatlas] Organ Seg. Chest & Abdomen Axial
8K
CTRATE [haaga2016ct] Report Gen. Chest Axial
26K
BIMCV-R [chen2024bimcv] Report Gen. Chest Axial
8K
ReXGroundingCT [baharoon2025rexgroundingct] Lesion Seg. Chest Axial 14 3K
Lesion Seg. Axial 340 10K
OmniAbnorm-CT-14K Lesion Det. Whole Body Coronal 255 2K
Report Gen. Sagittal 223 1.6K
Table 9: Comparison of key characteristics between OmniAbnorm-CT-14K and widely-used public CT imaging datasets. Note that CCC18 and DeepLesion has no category label for each annotated lesion. Even though some scans in these datasets are isotropic, they are acquired as axial-plane. Abbreviations: Category = Abnormality Category; Seg. = Segmentation; Det. = Detection; Gen = Generation.
E.4 Generation of Text Queries and Visual Prompts
@@
E.5 Detailed Distribution of OmniAbnorm-CT-14K
@@
Table 10: Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Table 11: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Ear Middle Ear Middle ear fluid or hemorrhage 5 0 0
@@
Pharyngeal space cystic mass 13 6 2
Pharyngeal emphysema 7 0 0
Others 5 7 1
Table 12: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Pharynx Pharyngeal space Pharyngeal space soft tissue mass 46 23 16
@@
Table 13: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
@@
Table 14: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Heart Myocardium Myocardial hypertrophy 4 1 1
@@
Coronary arteries Coronary artery myocardial bridge 1 0 0
@@
Table 15: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Table 16: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Gallbladder Gallbladder wall Gallbladder wall thickening 53 11 1
@@
Table 17: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
@@
Table 18: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Bladder Bladder cavity Bladder distention 4 0 1
Bladder stone 24 5 3
Bladder content density change 13 2 1
Others 4 3 0
@@
Testes, epididymis, and scrotum Testis Testicular enlargement 1 0 0
Testicular atrophy 0 0 0
Testicular soft tissue mass 1 0 0
Testicular calcification 0 0 0
Testicular cystic mass 0 0 0
Testicular torsion 0 0 0
Testicular hemorrhage and rupture 0 0 0
Others 0 0 0
Table 19: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Table 20: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
@@
Table 21: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
@@
Table 22: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.